OpenStudy (anonymous):

find the exact value: arctan(-1)

OpenStudy (experimentx):

find the value of \( \theta \) then \( \sin \theta = - \cos \theta \)

OpenStudy (anonymous):

Got it, thanks. One question though: wouldn't there be two answers for every problem of this type?

OpenStudy (anonymous):

I guess: \[\tan^{-1}(-x) \implies -\tan^{-1}(x)\]

OpenStudy (anonymous):

For example: the values of this one would be 7pi/4 and 3pi/4, correct?

OpenStudy (experimentx):

there would be many answers ... you have to restrict the domain. is there any restriction?

OpenStudy (anonymous):

No restriction, so the answer would be (3pi+4npi)/4, right?

OpenStudy (experimentx):

ok ok .. you can write it up then.

OpenStudy (anonymous):

Thanks!